# Gated Integration of Low-Rank Adaptation for Continual Learning of Language Models

### v1

@@ Abstract
Continual learning (CL), which requires the model to learn multiple tasks sequentially, is crucial for language models (LMs). Recently, low-rank adaptation (LoRA), one of the most representative parameter-efficient fine-tuning (PEFT) methods, has gained increasing attention in CL of LMs. However, most existing CL methods based on LoRA typically expand a new LoRA branch to learn each new task and force the new and old LoRA branches to contribute equally to old tasks, potentially leading to forgetting. In this work, we propose a new method, called gated integration of low-rank adaptation (GainLoRA), for CL of LMs. GainLoRA expands a new LoRA branch for each new task and introduces gating modules to integrate the new and old LoRA branches. Furthermore, GainLoRA leverages the new gating module to minimize the contribution from the new LoRA branch to old tasks, effectively mitigating forgetting and improving the model’s overall performance. Experimental results on CL benchmarks demonstrate that GainLoRA outperforms existing state-of-the-art methods. Code is available at https://github.com/liangyanshuo/gainlora.
@@ 1 Introduction
Continual learning (CL), which requires the model to learn multiple tasks sequentially, is crucial for language models (LMs) shi2024continual . Specifically, with extensive pre-trained knowledge and further fine-tuning strategies, existing LMs have demonstrated strong performance for a wide range of tasks brown2020language ; chowdhery2023palm ; touvron2023llama ; touvron2023llama2 ; zhang2022opt . However, when learning multiple tasks sequentially, LMs may lose knowledge acquired from old tasks, resulting in a significant degradation in performance on old tasks. This phenomenon, known as catastrophic forgetting luo2023empirical ; parisi2019continual ; DBLP:journals/pami/WangZSZ24 ; wang2023orthogonal , highlights the need for developing effective CL methods for LMs. Existing CL methods can be categorized into two main categories. The first category DBLP:conf/iclr/RazdaibiedinaMH23 assumes that task identities are available during inference, while the second category liang2024inflora ; zhao2024sapt tackles a more difficult and practical setting where task identities are unavailable during inference.
Recently, low-rank adaptation (LoRA) hu2021lora , one of the most representative parameter-efficient fine-tuning (PEFT) methods, has gained increasing attention in the CL of LMs bohaoscalable ; wang2023orthogonal . Specifically, by reparameterizing pre-trained weights in a low-rank form, LoRA updates only a limited number of parameters to adapt LMs to a downstream task, making the fine-tuning process much more efficient than updating all parameters of LMs han2024parameter . This efficiency also benefits CL, making LoRA increasingly popular in CL of LMs.
Most existing CL methods based on LoRA liang2024inflora ; zhao2024sapt typically expand a new LoRA branch for learning each new task while freezing all old LoRA branches. In this way, they avoid forgetting caused by directly updating the LoRA parameters of old tasks. However, to handle the practical CL scenario where task identities are unavailable at inference time, existing methods liang2024inflora ; smithcontinual ; wang2023orthogonal based on LoRA integrate new and old LoRA branches through a simple addition. Consequently, they force the new and old LoRA branches to contribute equally to old tasks, which means that the new LoRA branch may cause a relatively large change in the model’s output on old tasks. This leads to forgetting and degrades the model’s overall performance in CL.
In this work, we propose a new method, called gated integration of low-rank adaptation (GainLoRA), for CL of LMs. The contributions of GainLoRA are listed as follows:
@@
  • GainLoRA leverages the new gating module to minimize the contribution from the new LoRA branch to old tasks, effectively mitigating forgetting and improving the model’s overall performance.
@@ 2.1 Related Work
Parameter-Efficient Fine-Tuning Parameter-efficient fine-tuning (PEFT) methods tune a limited number of parameters to adapt a pre-trained model for downstream tasks, showing much more efficiency than tuning all the parameters of the pre-trained model, especially for LMs zhangadaptive . For example, Adapter houlsby2019parameter modifies the model architecture by introducing trainable modules into Transformer layers and tunes these modules for downstream tasks. Prompt-tuning DBLP:conf/emnlp/LesterAC21 and Prefix-tuning DBLP:conf/acl/LiL20 insert learnable tokens into the input and tune them for downstream tasks. Low-rank adaptation (LoRA) hu2021lora reparameterizes the original model parameters with low-rank matrices and tunes these matrices for downstream tasks. Although tuning significantly fewer parameters than full fine-tuning, PEFT can achieve comparable performance to full fine-tuning across a wide range of natural language processing (NLP) tasks fu2022adapterbias ; hu2021lora ; DBLP:conf/nips/MahabadiHR21 ; zaken2022bitfit .
@@
Many CL methods aranilearning ; li2017learning ; liang2023adaptive are designed to train models from scratch. Recent studies liang2024inflora ; DBLP:conf/iclr/RazdaibiedinaMH23 ; wang2023orthogonal ; wang2022learning have shown that leveraging PEFT strategies to fine-tune pre-trained models enables CL methods to achieve superior performance across tasks. Among various PEFT methods, LoRA has gained particular attention due to its strong performance and efficiency liang2024inflora ; smith2023construct ; wang2023orthogonal ; zhao2024sapt . However, most of these methods expand a new LoRA branch to handle each new task while freezing old LoRA branches to mitigate catastrophic forgetting. However, they force the new and old LoRA branches to contribute equally to old tasks, potentially leading to forgetting.
@@ 2.2 Preliminaries
Problem Definition We follow existing CL works wang2023orthogonal ; zhao2024sapt to formalize the problem definition for CL of LMs. Specifically, in CL, a sequence of tasks $\{\mathcal{T}_{1},\mathcal{T}_{2},...,\mathcal{T}_{T}\}$ is presented to the model sequentially, where $T$ denotes the total number of tasks. The $t$-th task $\mathcal{T}_{t}$ consists of a training dataset $\mathcal{D}_{t}$. For any given sample $(\bm{x}_{t},\bm{y}_{t})\in\mathcal{D}_{t}$, $\bm{x}_{t}$ denotes an input sentence and $\bm{y}_{t}$ denotes the corresponding output. When learning the $t$-th new task, the model is required to mitigate catastrophic forgetting of the $t-1$ previously learned tasks.
Similar to existing CL works for LMs bohaoscalable ; zhao2024sapt , we consider a more challenging CL setting defined by three key challenges: (1) the model is presented with a sequence of tasks spanning various types, such as dialogue generation, information extraction and so on; (2) the model is not provided with task identities at inference time; (3) the model must learn without access to real or synthetic samples from previously learned tasks.
Low-Rank Adaptation LoRA hu2021lora is a widely adopted PEFT method used for fine-tuning various pre-trained models, particularly LMs. Specifically, let $\bm{W}\in\mathbb{R}^{d_{out}\times d_{in}}$ represent a pre-trained weight in LMs, where $d_{in}$ and $d_{out}$ are the input and output dimensions, respectively. LoRA introduces an additional branch consisting of two matrices, $\bm{A}\in\mathbb{R}^{d_{out}\times r}$ and $\bm{B}\in\mathbb{R}^{r\times d_{in}}$, where $r\ll{\rm min}(d_{in},d_{out})$. LoRA then modifies the forward propagation of this layer as $\bm{e}=(\bm{W}+\bm{A}\bm{B})\bm{h}$. Here, $\bm{h}$ and $\bm{e}$ denote the input and output, respectively. $\bm{A}$ is initialized to $\bm{0}$, and $\bm{B}$ is initialized using a Gaussian distribution. During fine-tuning for downstream tasks, the pre-trained weight $\bm{W}$ remains frozen, and only the parameters $\bm{A}$ and $\bm{B}$ are fine-tuned.
@@ 3 Methodology
where $a_{i}$ is an integration coefficient that determines the contribution of the $i$-th LoRA branch to the input $\bm{h}$. Note that $\bm{W}_{t-1}$ is a zero matrix when $t=1$. As a result, the forward propagation in this layer is modified as
@@
Many existing CL methods based on LoRA liang2024inflora ; smith2023construct ; smithcontinual ; wang2023orthogonal ; zhao2024sapt share a similar architecture to our method, as illustrated in Figure 1 (a). However, these methods fix all coefficients $\{a_{i}\}_{i=1}^{t}$ in (1) to $1$, forcing the new and old LoRA branches to contribute equally to old tasks. As a result, the new LoRA branch introduces a change of $\bm{A}_{t}\bm{B}_{t}\bm{h}$ to the output for inputs $\bm{h}$ associated with old tasks, potentially leading to forgetting. Although some methods attempt to mitigate this forgetting by imposing regularization smithcontinual or orthogonality constraints liang2024inflora on the new LoRA branch, the fixed integration coefficients $\{a_{i}\}_{i=1}^{t}$ still limit their performance, as demonstrated by the experimental results presented in Section 4. Some method  zhao2024sapt does not force the new and old LoRA branches to contribute equally to old tasks but relies on replaying synthetic old samples to mitigate forgetting, making it unsuitable for the scenario considered in this work.
Different from existing methods, GainLoRA introduces an independent gating module $g_{i}(\cdot)$ for each task $\mathcal{T}_{i}$ to generate the integration coefficients ($1\leq i\leq T$). To mitigate the forgetting caused by the new task, GainLoRA leverages the gating module to minimize the contribution from the new LoRA branch to the old tasks. The details will be introduced in the following subsections.
@@ 3.1 Architecture of Gating Modules
$\bm{p}_{0}={\rm Pool}({\rm Token}(\bm{x})),~{}\bm{p}_{l}=\sigma(\bm{G}_{i,l}\bm{p}_{l-1}),~{}l\in\{1,2,...,L\},~{}g_{i}(\bm{x})=f(\bm{G}_{i,L+1}\bm{p}_{L}).$ (3)
Here, ${\rm Token}(\cdot)$ represents the tokenizer used in LMs to extract token embeddings from the input $\bm{x}$. ${\rm Pool}(\cdot)$ denotes an average pooling operation applied to the token embeddings to produce a fixed-size vector. $\sigma(\cdot)$ denotes the non-linear activation function. $\bm{G}_{i,l}$ denotes the weight matrix for the $l$-th layer of $g_{i}(\cdot)$ ($11\leq l\leq L+$ 1). In the final layer, $\bm{G}_{i,L+1}$ is a vector that maps the input vector $\bm{p}_{L+1}$ to a scalar. Following existing works with gating mechanisms cho2014properties ; DBLP:journals/neco/HochreiterS97 , the function $f(\cdot)$ is designed to map a scalar to a value within $[0,1]$, that is, $f(\cdot):\mathbb{R}\rightarrow[0,1]$.
Note that the input to gating modules is the same as that of LMs, denoted as $\bm{x}$, which differs from the input to LoRA in a specific layer, denoted as $\bm{h}$. During the learning of the $t$-th new task, only the new gating module $g_{t}(\cdot)$ is updated, while all the old gating modules $\{g_{i}(\cdot)\}_{i=1}^{t-1}$ remain frozen.
3.2 Minimizing the Contribution from the new LoRA branch to Old Tasks
GainLoRA minimizes the contribution from the new LoRA branch to old tasks by making $a_{t}=g_{t}(\bm{x})$ as close to 00 as possible for any input $\bm{x}$ from old tasks $\{\mathcal{T}_{i}\}_{i=1}^{t-1}$. However, since we focus on the scenario where no real or synthetic samples from old tasks are accessible, directly optimizing $g_{t}(\bm{x})$ to 00 is impractical. To overcome this challenge, GainLoRA imposes constraints on the new gating module $g_{t}(\cdot)$, implicitly guiding $g_{t}(\bm{x})$ to close to 00 and reduce the contribution of the new LoRA branch to old tasks.
In the following two subsections, we first describe the constraints imposed on the new gating module $g_{t}(\cdot)$ and explain how these constraints guide $g_{t}(\bm{x})$ close to 00 for any $\bm{x}$ from the old tasks. Then, we detail the implementation of these constraints during training.
@@ 3.2.1 Constraints on New Gating Module
To formalize the constraints imposed on the new gating module $g_{t}(\cdot)$, we define the subspace spanned by the inputs to $\bm{G}_{t,l}~{}(1\leq l\leq L+1)$ from the previous $t-1$ tasks as:
@@
where ${\rm Init}(\bm{G}_{t,L+1})$ denotes the initialization of $\bm{G}_{t,L+1}$. These constraints ensure that for any sample $\bm{x}$ from the old tasks, the integration coefficient satisfies $a_{t}=g_{t}(\bm{x})=f({\rm Init}(\bm{G}_{t,L+1})\bm{p}_{L})=0$, where $\bm{p}_{L}$ is defined in (3). The second equality holds since $\bm{G}_{t,L+1}={\rm Init}(\bm{G}_{t,L+1})$ before learning the $t$-th new task. The third equality holds because $f(0)=0$ and $\bm{p}_{L}\in\mathcal{M}_{t,L+1}$ for any $\bm{x}$ from previous $t-1$ tasks.
@@
If the constraints in (6) are satisfied, subspaces $\{\mathcal{M}_{t,l}\}_{l=1}^{L+1}$ remain unchanged during the learning of the $t$-th task. Furthermore, for any input $\bm{x}$ from the previous $t-1$ tasks, $g_{t}(\bm{x})$ remains unchanged during the learning of the $t$-th task.

### v2

# Gated Integration of Low-Rank Adaptation for Continual Learning of Large Language Models

@@ Abstract
Continual learning (CL), which requires the model to learn multiple tasks sequentially, is crucial for large language models (LLMs). Recently, low-rank adaptation (LoRA), one of the most representative parameter-efficient fine-tuning (PEFT) methods, has gained increasing attention in CL of LLMs. However, most existing CL methods based on LoRA typically expand a new LoRA branch to learn each new task and force the new and old LoRA branches to influence old tasks equally, potentially leading to forgetting. In this work, we propose a new method, called gated integration of low-rank adaptation (GainLoRA), for CL of LLMs. GainLoRA expands a new LoRA branch for each new task and introduces gating modules to integrate the new and old LoRA branches. Furthermore, GainLoRA leverages the new gating module to minimize the influence from the new LoRA branch to old tasks, effectively mitigating forgetting and improving the model’s overall performance. Experimental results on CL benchmarks demonstrate that GainLoRA outperforms existing state-of-the-art methods. Code is available at https://github.com/liangyanshuo/gainlora.
@@ 1 Introduction
Continual learning (CL), which requires the model to learn multiple tasks sequentially, is crucial for large language models (LLMs) shi2024continual . Specifically, although existing LLMs have demonstrated strong performance for a wide range of tasks brown2020language ; chowdhery2023palm ; touvron2023llama ; touvron2023llama2 ; zhang2022opt with extensive pre-trained knowledge and further fine-tuning strategies, they may lose knowledge acquired from old tasks when learning multiple tasks sequentially. This phenomenon, known as catastrophic forgetting luo2023empirical ; parisi2019continual ; DBLP:journals/pami/WangZSZ24 ; wang2023orthogonal , highlights the need for developing effective CL methods for LLMs. Existing CL methods can be categorized into two main categories. The first category DBLP:conf/iclr/RazdaibiedinaMH23 assumes that task identities are available during inference, while the second category liang2024inflora ; zhao2024sapt tackles a more difficult and practical setting where task identities are unavailable during inference.
Recently, low-rank adaptation (LoRA) hu2021lora , one of the most representative parameter-efficient fine-tuning (PEFT) methods, has gained increasing attention in the CL of LLMs bohaoscalable ; wang2023orthogonal . Specifically, by reparameterizing pre-trained weights in a low-rank form, LoRA updates only a limited number of parameters to adapt LLMs to a downstream task, making the fine-tuning process much more efficient than updating all parameters of LLMs han2024parameter . This efficiency also benefits CL, making LoRA increasingly popular in CL of LLMs.
Most existing CL methods based on LoRA liang2024inflora ; zhao2024sapt typically expand a new LoRA branch for learning each new task while freezing all old LoRA branches. In this way, they avoid forgetting caused by directly updating the LoRA parameters of old tasks. However, to handle the practical CL scenario where task identities are unavailable at inference time, existing methods liang2024inflora ; smithcontinual ; wang2023orthogonal based on LoRA integrate new and old LoRA branches through a simple addition. Consequently, they force the new and old LoRA branches to influence old tasks equally, which means that the new LoRA branch may cause a relatively large change in the model’s output on old tasks. This leads to forgetting and degrades the model’s overall performance in CL.
In this work, we propose a new method, called gated integration of low-rank adaptation (GainLoRA), for CL of LLMs. The contributions of GainLoRA are listed as follows:
@@
  • GainLoRA leverages the new gating module to minimize the influence from the new LoRA branch to old tasks, effectively mitigating forgetting and improving the model’s overall performance.
@@ 2.1 Related Work
Parameter-Efficient Fine-Tuning Parameter-efficient fine-tuning (PEFT) methods tune a limited number of parameters to adapt a pre-trained model for downstream tasks, showing much higher efficiency than tuning all the parameters of the pre-trained model, especially for LLMs zhangadaptive . For example, Adapter houlsby2019parameter modifies the model architecture by introducing trainable modules into Transformer layers and tunes these modules for downstream tasks. Prompt-tuning DBLP:conf/emnlp/LesterAC21 and Prefix-tuning DBLP:conf/acl/LiL20 insert learnable tokens into the input and tune them for downstream tasks. Low-rank adaptation (LoRA) hu2021lora reparameterizes the original model parameters with low-rank matrices and tunes these matrices for downstream tasks. Although tuning significantly fewer parameters than full fine-tuning, PEFT can achieve comparable performance to full fine-tuning across a wide range of natural language processing (NLP) tasks fu2022adapterbias ; hu2021lora ; DBLP:conf/nips/MahabadiHR21 ; zaken2022bitfit .
@@
Many CL methods aranilearning ; li2017learning ; liang2023adaptive are designed to train models from scratch. Recent studies liang2024inflora ; DBLP:conf/iclr/RazdaibiedinaMH23 ; wang2025self ; wang2023orthogonal ; wang2022learning have shown that leveraging PEFT strategies to fine-tune pre-trained models enables CL methods to achieve superior performance across tasks. For example, some methods DBLP:conf/iclr/QinJ22 ; DBLP:conf/iclr/RazdaibiedinaMH23 ; wang2022learning ; zhao2024sapt use prompt-tuning for continual learning. They either maintain independent prompts for each task or maintain a prompt pool and perform query-key matching to learn new tasks. There are also many methods liang2024inflora ; smith2023construct ; wang2023orthogonal ; zhao2024sapt adopting LoRA for continual learning. Most of these methods expand a new LoRA branch to handle each new task while freezing old LoRA branches to mitigate catastrophic forgetting. However, they force the new and old LoRA branches to influence old tasks equally, potentially leading to forgetting.
@@ 2.2 Preliminaries
Problem Definition We follow existing CL works wang2023orthogonal ; zhao2024sapt to formalize the problem definition for CL of LLMs. Specifically, in CL, a sequence of tasks $\{\mathcal{T}_{1},\mathcal{T}_{2},...,\mathcal{T}_{T}\}$ is presented to the model sequentially, where $T$ denotes the total number of tasks. The $t$-th task $\mathcal{T}_{t}$ consists of a training dataset $\mathcal{D}_{t}$. For any given sample $(\bm{x}_{t},\bm{y}_{t})\in\mathcal{D}_{t}$, $\bm{x}_{t}$ denotes an input sentence and $\bm{y}_{t}$ denotes the corresponding output. When learning the $t$-th new task, the model is required to mitigate catastrophic forgetting of the t1t-1 previously learned tasks.
Similar to existing CL works for LLMs bohaoscalable ; zhao2024sapt , we consider a more challenging CL setting with three key challenges: (1) the model is presented with a sequence of tasks spanning various types, such as dialogue generation, information extraction and so on; (2) the model is not provided with task identities at inference time; (3) the model must learn without access to real or synthetic samples from previously learned tasks.
Low-Rank Adaptation LoRA hu2021lora is a widely adopted PEFT method used for fine-tuning various pre-trained models, particularly LLMs. Specifically, let $\bm{W}\in\mathbb{R}^{d_{out}\times d_{in}}$ represent a pre-trained weight in LLMs, where $d_{in}$ and $d_{out}$ are the input and output dimensions, respectively. LoRA introduces an additional branch consisting of two matrices, $\bm{A}\in\mathbb{R}^{d_{out}\times r}$ and $\bm{B}\in\mathbb{R}^{r\times d_{in}}$, where $r\ll{\rm min}(d_{in},d_{out})$. LoRA then modifies the forward propagation of this layer as
$\bm{e}=(\bm{W}+\bm{A}\bm{B})\bm{h}.$
Here, $\bm{h}$ and $\bm{e}$ denote the input and output, respectively. $\bm{A}$ is initialized to $\bm{0}$, and $\bm{B}$ is initialized with a Gaussian distribution. During fine-tuning for downstream tasks, the pre-trained weight $\bm{W}$ remains frozen, and only the parameters $\bm{A}$ and $\bm{B}$ are fine-tuned.
@@ 3 Methodology
where $a_{i}$ is an integration coefficient that determines the influence of the $i$-th LoRA branch to the input $\bm{h}$. Note that $\bm{W}_{t-1}$ is a zero matrix when $t=1$. As a result, the forward propagation in this layer is modified as
@@
Many existing CL methods based on LoRA liang2024inflora ; smith2023construct ; smithcontinual ; wang2023orthogonal ; zhao2024sapt share a similar architecture to our method, as illustrated in Figure 1 (a). However, these methods fix all coefficients $\{a_{i}\}_{i=1}^{t}$ in (1) to $1$, forcing the new and old LoRA branches influence old tasks equally. As a result, the new LoRA branch introduces a change of $\bm{A}_{t}\bm{B}_{t}\bm{h}$ to the output for inputs $\bm{h}$ associated with old tasks, potentially leading to forgetting. Although some methods attempt to mitigate this forgetting by imposing regularization smithcontinual or orthogonality constraints liang2024inflora on the new LoRA branch, the fixed integration coefficients $\{a_{i}\}_{i=1}^{t}$ still limit their performance, as demonstrated by the experimental results presented in Section 4. The method in zhao2024sapt does not force the new and old LoRA branches to influence old tasks equally but relies on replaying synthetic old samples to mitigate forgetting, making it unsuitable for the scenario considered in this work.
Different from existing methods, GainLoRA introduces an independent gating module $g_{i}(\cdot)$ for each task $\mathcal{T}_{i}$ to generate the integration coefficients ($1\leq i\leq T$). To mitigate the forgetting caused by the new task, GainLoRA leverages the gating module to minimize the influence from the new LoRA branch to the old tasks. The details will be introduced in the following subsections.
@@ 3.1 Architecture of Gating Modules
$\bm{p}_{i,0}$ $=\bm{p}_{0}={\rm Pool}({\rm Token}(\bm{x})),$
$\bm{p}_{i,l}$ $=\sigma(\bm{G}_{i,l}\bm{p}_{i,l-1}),~l\in\{1,2,...,L\},$
$g_{i}(\bm{x})$ $=f(\bm{G}_{i,L+1}\bm{p}_{i,L}).$ (3)
Here, ${\rm Token}(\cdot)$ represents the tokenizer used in LLMs to extract token embeddings from the input $\bm{x}$. ${\rm Pool}(\cdot)$ denotes an average pooling operation applied to the token embeddings to produce a fixed-size vector. $\sigma(\cdot)$ denotes the non-linear activation function. $\bm{G}_{i,l}$ denotes the weight matrix for the $l$-th layer of $g_{i}(\cdot)$ ($1\leq l\leq L+1$). In the final layer, $\bm{G}_{i,L+1}$ is a vector that maps the input vector $\bm{p}_{i,L}$ to a scalar. Following existing works with gating mechanisms cho2014properties ; DBLP:journals/neco/HochreiterS97 , the function $f(\cdot)$ is designed to map a scalar to a value within $[0,1]$, that is, $f(\cdot):\mathbb{R}\rightarrow[0,1]$.
Note that the input to gating modules is the same as that of LLMs, denoted as $\bm{x}$, which differs from the input to LoRA in a specific layer, denoted as $\bm{h}$. During the learning of the $t$-th new task, only the new gating module $g_{t}(\cdot)$ is updated, while all the old gating modules $\{g_{i}(\cdot)\}_{i=1}^{t-1}$ remain frozen.
3.2 Minimizing the Influence from the New LoRA Branch to Old Tasks
GainLoRA minimizes the influence from the new LoRA branch to old tasks by making $a_{t}=g_{t}(\bm{x})$ as close to 0 as possible for any input $\bm{x}$ from old tasks $\{\mathcal{T}_{i}\}_{i=1}^{t-1}$. However, since we focus on the scenario where no real or synthetic samples from old tasks are accessible, directly optimizing $g_{t}(\bm{x})$ to 0 is impractical. To overcome this challenge, GainLoRA imposes constraints on the new gating module $g_{t}(\cdot)$, implicitly guiding $g_{t}(\bm{x})$ to close to 0 and reduce the influence of the new LoRA branch to old tasks.
In the following two subsections, we first describe the constraints imposed on the new gating module $g_{t}(\cdot)$ and explain how these constraints guide $g_{t}(\bm{x})$ close to 0 for any $\bm{x}$ from the old tasks. Then, we detail the implementation of these constraints during training.
@@ 3.2.1 Constraints on New Gating Module
To formalize the constraints imposed on the new gating module $g_{t}(\cdot)$, we define the subspace spanned by the inputs to $\bm{G}_{t,l}~(1\leq l\leq L+1)$ from the previous t1t-1 tasks as:
@@
where ${\rm Init}(\bm{G}_{t,L+1})$ denotes the initialization of $\bm{G}_{t,L+1}$. These constraints ensure that for any sample $\bm{x}$ from the old tasks, the integration coefficient satisfies
$a_{t}=g_{t}(\bm{x})=f({\rm Init}(\bm{G}_{t,L+1})\bm{p}_{t,L})=0,$ (6)
where $\bm{p}_{t,L}$ is defined in (3.1). The second equality holds since $\bm{G}_{t,L+1}={\rm Init}(\bm{G}_{t,L+1})$ before learning the $t$-th new task. The third equality holds because $f(0)=0$ and $\bm{p}_{t,L}\in\mathcal{M}_{t,L+1}$ for any $\bm{x}$ from previous t1t-1 tasks.
@@
If the constraints in (7) are satisfied, subspaces $\{\mathcal{M}_{t,l}\}_{l=1}^{L+1}$ remain unchanged during the learning of the $t$-th task. Furthermore, for any input $\bm{x}$ from the previous t1t-1 tasks, $g_{t}(\bm{x})$ remains unchanged during the learning of the $t$-th task.